OpenStudy (lovelycharm):

solve each system by graphing. tell weather the system has one solution, infinitely many solution, or one solution y = x - 1 y = -3x - 5

OpenStudy (zpupster):

first graph

OpenStudy (zpupster):

where they intersect is our solution

OpenStudy (zpupster):

confirming we can do something like solving by multiplying the 2nd equation by (-1) and then add the 2 equations together. 0= 4x+4 -4=4x divide both sides by 4 getting x by itself -1=x plugging in y=(-1)-1 y=-2 checking y=-3(-1) -5 y=3-5 y=-2

OpenStudy (lovelycharm):

so the answer is one solution @zpupster

OpenStudy (lovelycharm):

where are u why u dont answer @zpupster